.In a right triangle, the 58-cm hypotenuse makes a 51-degree angle with one of the legs. To the nearest tenth of a cm, how long is that leg? Once you have the answer, find two ways to calculate the length of the other leg. They should both give the same answer.

Answer by Alan3354(69443)

.In a right triangle, the 58-cm hypotenuse makes a 51-degree angle with one of the legs. To the nearest tenth of a cm, how long is that leg?
------
Call the leg a
a = 58*cos(51)
===================

Once you have the answer, find two ways to calculate the length of the other leg. They should both give the same answer.
------------
They should both give the same answer. --- as opposed to only 1 giving the same answer?
---
Try "They should give the same answer."
---------------------------
The other side (call it b) is 58*sin(51)
---
And b^2 = 58^2 - a^2